MAGNETIC HYPERFINE INTERACTIONS IN ^{171}YbF AND ^{173}YbF^{a}.

\underline{TIMOTHY C. STEIMLE}, TONGMEI MA, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ, USA 85287; COLAN LINTON, Physics Department, University of New Brunswick, Fredericton, NB, Canada E3B 5A3.

The heavy polar molecule ytterbium monofluoride, YbF, has been used to set an upper limit on the electric-dipole moment of the electron, d_{e}^{b}. The extraction of d_{e} from experimental data requires knowledge of the effective electric field in the region of the Yb nucleus, which can only be obtained from electron structure calculations^{c}. The most effective means of gauging the quality of the predicted wavefunction in the region of the nucleus is via a comparison with experimental measurements of the magnetic hyperfine interactions. Here we report on ^{171}Yb(I=1/2), ^{173}Yb(I=5/2), and 19F(I=1/2) magnetic hyperfine interactions from the analysis of the A^{2}Π_{1/2}−X^{2}Σ^{+} (0,0) band of ytterbium monofluoride, YbF, recorded at the near-natural linewidth limit. The results are compared with ab initio predictions^{b}, the microwave measurements for ^{174}YbF^{d}, and the matrix-isolated ESR results for ^{171}YbF^{e}.

aSupported by NSF-Experimental Physical Chemistry

bJ.J. Hudson, B.E. Sauer, M.R. Tarbutt, and E.A. Hinds, Phys. Rev. Lett. 98, 023003-4 (2002).

cA.V. Titov, N.S. Mosyagin, A.N. Petrov, and T.A. Iseav, Int. J. Quantum Chem. 104, 223-239 (2005).

dC.S. Dickinson, J.A. Coxon, N.R. Walker, and M.C.L. Gerry, J. Chem. Phys. 115, 6979-6989 (2001).

eR.J. Van Zee, M.L. Seely, T.C. DeVore, and W. Weltner, Jr. J. Chem. Phys. 82, 1192-1194 (1978).